# Formal extension of noncommutative tensor-triangular support varieties

### v2

@@ 1. Introduction
Since the early 80s, support varieties have been an important tool in representation theory, when cohomological support and rank support were introduced to provide geometric techniques for studying categories of algebraic origin [AS82, Car83, Alp87]; analogous versions of support in commutative algebra and algebraic geometry were being developed around the same period [Hop87, Avr89, Nee92a]. A support theory here and throughout means a map which assigns to each object of some category a closed subset of some topological space, in a way which is compatible with the structure of the category. By the early 2000s, rank varieties were generalized to finite group schemes via the $\pi$-points of Friedlander–Pevtsova [FP07]. With the advent of tensor-triangular geometry, a new support theory (the Balmer support, universal support, or tensor-triangular support) based on the Balmer spectrum entered the literature [Bal05]. It has found widespread applications in algebraic topology, commutative algebra, algebraic geometry, modular representation theory, equivariant stable homotopy theory, and motivic homotopy theory (see [Bal20] and the references therein). Noncommutative Balmer spectra were developed in [NVY22a] to study monoidal triangulated categories without braidings, which arise naturally from various settings such as stable categories of finite tensor categories (e.g.  the representation theory of finite-dimensional Hopf algebras), group actions on tensor-triangulated categories via the crossed product (see e.g. [HV]), and derived categories of bimodules over a noncommutative ring (see [BKSS20, Section 1]).
@@
The paper is organized as follows: in Section 2 we review background on monoidal triangular geometry and support varieties. In Section 3, we define extended support via Rickard idempotents in the noncommutative case, based on the classical commutative constructions. In Section 4, we prove that the extension of the Balmer support satisfies a generalized tensor prodouct property, and along the way give some useful descriptions of the extended Balmer supports of the Rickard idempotents themselves. In Section 5, we prove that any extension of a support equipped with a comparison map (that is, a map of topological spaces with compatibility properties that allow a “comparison” between the support in question and the Balmer support) is faithful, as long as either the Balmer spectrum of $\mathbf{K}^{c}$ is Noetherian or $\mathbf{K}^{c}$ is has a thick generator. In Section 6, we show that if the Balmer spectrum is Noetherian, then any faithful and realizing support which satisfies a generalized tensor product property and admits a faithful extension also admits a comparison map. In Section 7, we prove that if a support map is induced by a surjective map from a Noetherian Balmer spectrum, then the extension of that support is faithful. Lastly, in Section 8, we turn our attention to the specific case of stable categories of finite tensor categories, and prove Theorem 1.2, that is, that extended central cohomological support is faithful under certain assumptions.
@@ 2. Monoidal triangular geometry and support theories
See [Bal05] for the case that $\mathbf{K}^{c}$ is commutative, [MR23] for the case that $\mathbf{K}^{c}$ is completely prime, [Row] for the case that $\mathbf{K}^{c}$ is $\operatorname{Spc}$-Noetherian, and [NVY] for the case that $\mathbf{K}^{c}$ is finitely-generated (i.e., has a thick generator). ∎
@@ 3. Rickard idempotents and formal extensions of support
(1) is clear. (2) and (3) follow from [NVY, Lemma 3.2.1]. (4) is a direct consequence of (1), (2), and the fact that ${\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\in\operatorname{Loc}(\mathbf{I})$. (5) is straightforward. For (6), see [BF11, Remark 2.9].∎
@@ 5. Faithfulness of extended comparative support
by Lemma 3.2. But $L_{V}L_{W}A\in\langle\Theta(V)\cup\Theta(W)\rangle^{\perp}$ for any object $A$, by [NVY, Lemma 3.2.2(a)]. Now note that $L_{V}L_{W}{\it{\Gamma}}_{V\cup W}\mathbf{1}\in\Theta(V\cup W)=\langle\Theta(V)\cup\Theta(W)\rangle$ by Lemma 5.5, and it follows that $L_{V}L_{W}{\it{\Gamma}}_{V\cup W}\mathbf{1}\cong 0$, hence $L_{V}L_{W}\mathbf{1}\cong L_{V\cup W}\mathbf{1}$. ∎
@@ 8. Applications to stable categories of finite tensor categories
where $G$ is the sum of simple objects. So long as $\mathbf{I}$ and $\mathbf{M}$ are disjoint, we may produce a prime ideal containing $\mathbf{I}$ which intersects $\mathbf{M}$ trivially, by [NVY, Lemma A.1.1], since $\mathbf{M}$ has the property that given any two objects $B$ and $D$, there exists an object $C$ such that $B\otimes C\otimes D\in\mathbf{M}$ (of course, we have constructed it so that $D$ can be chosen as $G$). Indeed, $\mathbf{I}$ and $\mathbf{M}$ must be disjoint, since by [NVY24, Proposition 7.1.1(d)] we have $\operatorname{supp}^{\rm C}(\operatorname{cone}(g))=Z(g)$ for any $g\in C^{\bullet}$, hence
@@
for any $B\in\mathbf{I}$. Therefore, $\mathfrak{p}\not\in\operatorname{supp}^{\rm C}(B)$ for any $B\in\mathbf{I}$. Since $\operatorname{supp}^{\rm C}(A\otimes G\otimes A\otimes...\otimes G\otimes A)=\operatorname{supp}^{\rm C}(A)$ by Proposition 8.3, and $\mathfrak{p}\in\operatorname{supp}^{\rm C}(A)$, it follows that $\mathbf{I}\cap\mathbf{M}$ is empty. Thus, by [NVY, Lemma A.1.1], we can produce a prime ideal $\mathbf{P}$ containing $\mathbf{I}$ and disjoint from $\mathbf{M}$. Since $\mathbf{P}$ contains $\mathbf{I}$, we have

### v3

@@ 1. Introduction
Since the early 80s, support varieties have been an important tool in representation theory, when cohomological support and rank support were introduced to provide geometric techniques for studying categories of algebraic origin [AS82, Car83, Alp87]; analogous versions of support in commutative algebra and algebraic geometry were being developed around the same period [Hop87, Avr89, Nee92a]. A support theory here and throughout means a map which assigns to each object of some category a closed subset of some topological space, in a way which is compatible with the structure of the category. By the early 2000s, rank varieties were generalized to finite group schemes via the $\pi$-points of Friedlander–Pevtsova [FP07]. With the advent of tensor-triangular geometry, a new support theory (the Balmer support, universal support, or tensor-triangular support) based on the Balmer spectrum entered the literature [Bal05]. It has found widespread applications in algebraic topology, commutative algebra, algebraic geometry, modular representation theory, equivariant stable homotopy theory, and motivic homotopy theory (see [Bal20] and the references therein). Noncommutative Balmer spectra were developed in [NVY22a] to study monoidal triangulated categories without braidings, which arise naturally from various settings such as stable categories of finite tensor categories (e.g.  the representation theory of finite-dimensional Hopf algebras), group actions on tensor-triangulated categories via the crossed product (see e.g. [HV23]), and derived categories of bimodules over a noncommutative ring (see [BKSS20, Section 1]).
@@
The paper is organized as follows: in Section 2 we review background on monoidal triangular geometry and support varieties. In Section 3, we define extended support via Rickard idempotents in the noncommutative case, based on the classical commutative constructions. In Section 4, we prove that the extension of the Balmer support satisfies a generalized tensor prodouct property, and along the way give some useful descriptions of the extended Balmer supports of the Rickard idempotents themselves. In Section 5, we prove that any extension of a support equipped with a comparison map (that is, a map of topological spaces with compatibility properties that allow a “comparison” between the support in question and the Balmer support) is faithful, as long as either the Balmer spectrum of $\mathbf{K}^{c}$ is Noetherian or $\mathbf{K}^{c}$ is has a thick generator. In Section 6, we show that if the Balmer spectrum is Noetherian, then any faithful and realizing support which satisfies a generalized tensor product property and admits a faithful extension also admits a comparison map. In Section 7, we prove that if a support map is induced by a surjective map from a Noetherian Balmer spectrum, then the extension of that support is faithful. Lastly, in Section 8, we turn our attention to the specific case of stable categories of finite tensor categories, and prove Theorem 1.2, that is, that extended central cohomological support is faithful under certain assumptions. We then apply these results to answer a question of Pevtsova–Witherspoon [PW15], which recovers using different techniques a theorem of Negron–Pevtsova [NP23a].
Remark 1.3 (Note on further applications).
Beyond the applications to finite tensor categories presented in Section 8, the results of this paper are applied in joint work of the second author with D.K.  Nakano and M.T.  Yakimov [NVY25] to study stratification in monoidal triangular geometry and prove a version of Balmer’s Nerves of Steel Conjecture for crossed product categories. Future applications will also include work currently in preparation joint between the second author with Ø.  Solberg and S.  Witherspoon to study Balmer spectra of some monoidal triangulated categories related to bimodules for self-injective algebras.
@@ 2. Monoidal triangular geometry and support theories
See [Bal05] for the case that $\mathbf{K}^{c}$ is commutative, [MR23] for the case that $\mathbf{K}^{c}$ is completely prime, [Row24] for the case that $\mathbf{K}^{c}$ is $\operatorname{Spc}$-Noetherian, and [NVY23] for the case that $\mathbf{K}^{c}$ is finitely-generated (i.e., has a thick generator). ∎
@@ 3. Rickard idempotents and formal extensions of support
(1) is clear. (2) and (3) follow from [NVY23, Lemma 3.2.1]. (4) is a direct consequence of (1), (2), and the fact that ${\it{\Gamma}}_{\mathbf{I}}\mathbf{1}\in\operatorname{Loc}(\mathbf{I})$. (5) is straightforward. For (6), see [BF11, Remark 2.9].∎
@@ 5. Faithfulness of extended comparative support
by Lemma 3.2. But $L_{V}L_{W}A\in\langle\Theta(V)\cup\Theta(W)\rangle^{\perp}$ for any object $A$, by [NVY23, Lemma 3.2.2(a)]. Now note that $L_{V}L_{W}{\it{\Gamma}}_{V\cup W}\mathbf{1}\in\Theta(V\cup W)=\langle\Theta(V)\cup\Theta(W)\rangle$ by Lemma 5.5, and it follows that $L_{V}L_{W}{\it{\Gamma}}_{V\cup W}\mathbf{1}\cong 0$, hence $L_{V}L_{W}\mathbf{1}\cong L_{V\cup W}\mathbf{1}$. ∎
@@ 8. Applications to stable categories of finite tensor categories
where $G$ is the sum of simple objects. So long as $\mathbf{I}$ and $\mathbf{M}$ are disjoint, we may produce a prime ideal containing $\mathbf{I}$ which intersects $\mathbf{M}$ trivially, by [NVY23, Lemma A.1.1], since $\mathbf{M}$ has the property that given any two objects $B$ and $D$, there exists an object $C$ such that $B\otimes C\otimes D\in\mathbf{M}$ (of course, we have constructed it so that $D$ can be chosen as $G$). Indeed, $\mathbf{I}$ and $\mathbf{M}$ must be disjoint, since by [NVY24, Proposition 7.1.1(d)] we have $\operatorname{supp}^{\rm C}(\operatorname{cone}(g))=Z(g)$ for any $g\in C^{\bullet}$, hence
@@
for any $B\in\mathbf{I}$. Therefore, $\mathfrak{p}\not\in\operatorname{supp}^{\rm C}(B)$ for any $B\in\mathbf{I}$. Since $\operatorname{supp}^{\rm C}(A\otimes G\otimes A\otimes...\otimes G\otimes A)=\operatorname{supp}^{\rm C}(A)$ by Proposition 8.3, and $\mathfrak{p}\in\operatorname{supp}^{\rm C}(A)$, it follows that $\mathbf{I}\cap\mathbf{M}$ is empty. Thus, by [NVY23, Lemma A.1.1], we can produce a prime ideal $\mathbf{P}$ containing $\mathbf{I}$ and disjoint from $\mathbf{M}$. Since $\mathbf{P}$ contains $\mathbf{I}$, we have
@@
Example 8.8.
[…]
Pevtsova and Witherspoon prove in [PW15, Theorem 1.1] that the cohomological support of $\mathbf{K}^{c}$ is tensorial, and hence we have a continuous map
[…]
$𝕜\rho:\operatorname{Spc}\mathbf{K}^{c}\to\operatorname{Proj}\operatorname{H}^{\bullet}(A,\Bbbk)$
recalled above is a left inverse to $\eta$, and is closed by Proposition 8.4. By the classification of thick ideals for $\mathbf{K}^{c}$, also due to Pevtsova–Witherspoon [PW15, Theorem 1.2], $\rho$ must be injective, and hence $\eta$ and $\rho$ are inverse homeomorphisms. Indeed, using the explict formula for $\rho$, if $\rho(\mathbf{P})=\rho(\mathbf{Q})$ for $\mathbf{P},\mathbf{Q}\in\operatorname{Spc}\mathbf{K}^{c}$, then it would follow that for all $𝕜g\in\operatorname{H}^{\bullet}(A,\Bbbk)$, we would have $\operatorname{cone}(g)\in\mathbf{P}\Leftrightarrow\operatorname{cone}(g)\in\mathbf{Q}$, which would imply that the supports of $\mathbf{P}$ and $\mathbf{Q}$ would be the same, by [NVY24, Proposition 7.1.1], a contradiction. Therefore
[…]
These results allow us to apply Theorem 8.6, since $\operatorname{Proj}C^{\bullet}$ is Noetherian, $\operatorname{supp}^{\rm C}$ is tensorial, $C^{\bullet}$ is finitely-generated, and $\operatorname{Spc}\mathbf{K}^{c}$ is Noetherian. We obtain that $\operatorname{supp}^{\rm C}$ admits a faithful extension to all of $\mathbf{K}$. Furthermore, since $\operatorname{supp}^{\rm C}$ identifies with the Balmer support, by Proposition 4.4 we have tensoriality of $\operatorname{Supp}^{\rm C}$: the question posed by Pevtsova–Witherspoon, asking whether there exists an extension of $\operatorname{supp}^{\rm C}$ satisfying a tensor product property, has a positive answer.
We note that this problem was first solved in [NP23a] using an alternate method. Negron–Pevtsova construct a different extended support, termed hypersurface support, and prove in [NP23a, Theorem 7.9] that this hypersurface support is a faithful extended support satisfying a version of the tensor product property. However, for hypersurface support to be defined, one needs to have a smoothly integrable Hopf algebra. We emphasize that the techniques we use to recover this result above make no reference to the smooth integration, and that by Proposition 4.4 and Corollary 5.11 we expect existence of a faithful and tensorial extended support to be a general phenomena in monoidal triangular geometry, and not specific to smoothly integrable Hopf algebras.